Volcán Barva (01.01.2017 to 03.31.2017)

You are here

PDF versionPDF version
Year: 
2017
Quarter: 
January - March
Communications activities: 

On February, the site manager attended a meeting of researchers working with camera traps in Costa Rica in order to do efforts in conservation together.

On February, the site manager attended using a TEAM stand, the “day of the water” activity organized by local people in Sarapiqui region.

On March, the Site Manager gave a TEAM talk to students of REU (Research Experience for Undergraduates) in OTS.

On March, the site manager attended a workshop to share TEAM Jaguar’s data in order to do a national analysis of actual distribution and occupancy rates of the species. This initiative is coordinated by SINAC (National System of Conservation Areas in Costa Rica).

On March, the site manager attended using a TEAM stand, the “multicultural day” activity organized by EARTH University Guacimo, CR.

Unusual Events at the TEAM Site: 

There were not unusual events in this sampling period.

New Species: 

There were not new species in this sampling period.

Protocol Activities: 

The annual census of trees, palms and lianas ≥10 cm diameter in the 1 ha plot located at 50, 100, 200, 400, 960, and 2800m asl was completed up to date. More than 4000 stems have been measured.

We have completed the terrestrial vertebrates monitoring for the sampling period in array 1 and 2 (low and middle elevation). 40 camera traps were deployed. Interesting sightings of 18 mammals were recorded included felids such as puma and ocelots were observed. Currently we are proccessing the images from array 2. 1500 images were processed on the deskteam for Array 1. All data (tpk file) were uploaded in the TEAm server.

Protocol Problems: 

We have several problems with Climate Protocol implementation. Since last year we have serious problems with the weather station possibly by humidity. At this moment the weather is not working and we need urgently to hire an expert to fix it. We are schedulling a visit to the weather station in two weeks with a person who is going to check it.

The other problem is that the access to reach the weather station located at middle elevation is difficult now because one of the bridges collapsed and it needed to walk around three hours to go there.

 

Schedule Problems: 

We have a small delay in the schedule but all sampling protocols will be finished on time. This is because we start the hire contract with OTS two weeks later to the schedule plan.

Logistical Problems: 

There is a bridge collapsed to go to middle elevation and to access to the weather station, and two vegetation plots (600, 960 masl). So the time to go there walking is very hard.

Suggestions: 

There is no suggestions to this sampling period